OpenStudy (anonymous):

Solve. 1. z/7 - 6 = 3 2. 3 - 2k = -7

OpenStudy (anonymous):

\[\frac{ z }{7}-6=3\]

OpenStudy (anonymous):

1st add 6 to both sides what do you get?

OpenStudy (anonymous):

16 ?

OpenStudy (anonymous):

6 +3 = ?

OpenStudy (anonymous):

9

OpenStudy (anonymous):

ok soo you get \[\frac{ z }{7 }=9\]

OpenStudy (anonymous):

Thank you, but I need help on number 2 also. Because it's a different equation.

OpenStudy (anonymous):

because z/7 -6=3 +6 +6 ok that was the first step

OpenStudy (anonymous):

now multyply 7 to both sides

OpenStudy (anonymous):

14

OpenStudy (anonymous):

\[(\frac{ 7 }{1 })\frac{ z }{ 7}=9(\frac{ 7 }{ 1 })\]

OpenStudy (anonymous):